Using X as a variable, translate twice a number increased by 7 equals 19 solve for X
Answer by Maths68(1474)

Let number = x
Twice a number = 2x
Increased by 7 = 2x+7
2x+7=19
2x=19-7
2x=12
2x/2=12/2
x=6
Number is = x = 6